# Dynamic Degradation Decomposition Network for All-in-One Image Restoration

### v1

@@ Abstract
Currently, restoring clean images from a variety of degradation types using a single model is still a challenging task. Existing all-in-one image restoration approaches struggle with addressing complex and ambiguously defined degradation types. In this paper, we introduce a dynamic degradation decomposition network for all-in-one image restoration, named D3Net. D3Net achieves degradation-adaptive image restoration with guided prompt through cross-domain interaction and dynamic degradation decomposition. Concretely, in D3Net, the proposed Cross-Domain Degradation Analyzer (CDDA) engages in deep interaction between frequency domain degradation characteristics and spatial domain image features to identify and model variations of different degradation types on the image manifold, generating degradation correction prompt and strategy prompt, which guide the following decomposition process. Furthermore, the prompt-based Dynamic Decomposition Mechanism (DDM) for progressive degradation decomposition, that encourages the network to adaptively select restoration strategies utilizing the two-level prompt generated by CDDA. Thanks to the synergistic cooperation between CDDA and DDM, D3Net achieves superior flexibility and scalability in handling unknown degradation, while effectively reducing unnecessary computational overhead. Extensive experiments on multiple image restoration tasks demonstrate that D3Net significantly outperforms the state-of-the-art approaches, especially improving PSNR by 5.47dB and 3.30dB on the SOTS-Outdoor and GoPro datasets, respectively.
@@
Image restoration, Prompt learning, Dynamic network.
@@ I Introduction
Figure 1: Conceptual differences between our work and the three main existing all-in-one methods (better viewed by zooming in).
Image restoration aims to remove degradations from corrupted images and generate high-quality ones. It is widely applied in various real-world scenarios, such as autonomous driving, digital photography, and security surveillance. Currently, most image restoration approaches focus on overcoming a single type of degradations, which may be denoising [1, 2, 3, 4, 5], deblurring [6, 7, 8, 9, 10], deraining [11, 12, 13, 14, 15], dehazing [16, 17, 18, 19], or low-light enhancement [20, 21, 21, 22, 23, 24, 25]. Although these previous research has yielded promising results in a specific application, the challenge of addressing multiple degradation tasks in image restoration with a single model still persists. This significantly hampers the practical applicability of the existing image restoration techniques.
@@
As shown in Fig. 1(d), in this paper, we propose the Dynamic Degradation Decomposition Network (D3Net) for all-in-one image restoration, which aims to address unknown multiple degradation situations. Unlike previous all-in-one image restoration approaches [26, 27, 28, 29, 30, 31], our method departs from the perspective of frequency domain and dynamic networks to effectively handle unknown and complex degradations without relying on explicit degradation labels or predefined prompts. In order to address the first question as above, we design a Cross-Domain Degradation Analyzer (CDDA) that generates degradation-related prompt to guide the degradation decomposition process. Specifically, it considers the complementary degradation characteristics between frequency and spatial domains, deeply interacting key degradation features in the frequency domain with spatial image features to identify and model the degradation types on the image manifold. Meanwhile, it generates degradation correction prompt and strategy prompt, thereby providing guidance for the following degradation decomposition process. Moreover, for the second question, we introduce a degradation-aware Dynamic Decomposition Mechanism (DDM). The design of DDM aims to progressively and dynamically decompose degradation features, fully leveraging the two-level prompt generated by CDDA to guide the image restoration process. Specifically, this mechanism distinctly enables the network to dynamically orchestrate and calibrate the input from diverse decomposition blocks, tailoring its model to suit the different degradation scenarios.
@@
  • We further introduce a novel dynamic decomposition mechanism. Owing to its adaptive prompt-guided mode, this mechanism can dynamically select an appropriate restoration strategy based on the degradation state of the images, thereby achieving significant degradation decomposition effects while reducing computational overhead.
@@ II-A Image Restoration
Image Restoration for Single Degradation. Traditional image restoration techniques mainly target at a specific type of degradation, such as blur, noise, or haze. These approaches [32, 33, 34, 35, 36], optimized for specific problems, can often achieve good results in the corresponding degradation scenario. For example, Mohan et al. [32] propose a model-based approach for addressing blur caused by camera shake. However, such approaches typically rely on specific prior degradation models, such as linear blur or motion blur. These model-based approaches may be severely limited when the actual degradation process does not match the assumed one. Recently, with the advancement of deep learning, numerous image restoration approaches based on deep learning have emerged and performed excellently across various image restoration tasks, such as denoising [1, 2, 3, 4, 5], deblurring [6, 7, 8, 9, 10], deraining [11, 12, 13, 14, 15], dehazing [16, 17, 18, 19], and low-light enhancement [20, 21, 21, 22, 23, 24, 25]. However, constrained by model architectures designed for single tasks, they tend to perform exceptionally well in single degradation removal tasks but struggle to deal with other types of degradation.
@@
Similarly, Conde et al. [31] propose a method utilizing predefined degradation texts to guide multiple degradation types. However, this approach does not deeply model the inherent degradation features within the image. By focusing solely on user-provided textual instructions without in-depth analysis of degradation patterns in the image data, it may overlook critical details necessary for optimal restoration, especially when degradations are subtle or intertwined.
@@ II-B Dynamic Networks
The research on dynamic networks has gradually garnered attention in the academic community, with the objective of reducing computational complexity while ensuring model performance. In recent years, researchers have introduced various methodologies to achieve this goal. For example,  [40] introduces the groundbreaking BL-DNN approach, which concatenates two distinct deep CNNs with varying depths. When the softmax prediction accuracy of the first CNN surpasses a predefined threshold, the network terminates prematurely to reduce computational complexity. Several subsequent works [41, 42, 43, 44] have also improved recognition tasks based on this concept. Moreover, some researchers develop methods that dynamically adjust convolutional kernel weights [45, 46, 47, 48] and kernel shapes [49, 50, 51] during inference based on their inputs. However, due to their irregular memory access and computation patterns, these methods are limited in further efficiency improvement. Diverging from most existing dynamic networks, we design a degradation prompt-based dynamic decomposition mechanism that dynamically selects appropriate image restoration strategies based on the state of image degradation.
@@ III-A The Architecture of D3Net
In order to achieve prompt-guided adaptive image degradation restoration, we design a D3Net as illustrated in Fig. 2. It primarily consists of a restoration reconstruction branch and a degradation decomposition branch. The restoration reconstruction branch employs a generic U-Net structure [52] to reconstruct the input image. Building upon this foundational process, for a given degraded input image $I$, the shallow degradation features $x$ extracted by the first encoder of the restoration reconstruction branch are inputted into the degradation decomposition branch. That is, $x=f_{\text{encoder}}(I)$. Concurrently, $I$ is fed into the cross-domain degradation analyzer, which generates degradation correction prompt $C_{t}$ and strategy prompt $S_{t}$ to guide the degradation decomposition branch. The dynamic decomposition branch selects degradation-adaptive strategies under the guidance of $x$, $C_{t}$, and $S_{t}$ to decompose $x$ into degradation correct features $Y$ and restoration strategy $x_{d}$, i.e. $Y,x_{d}=f_{\text{dec}}(x,C_{t},S_{t})$. Further, $Y$ generates the degradation-decomposed features $I^{\prime}$ under the guidance of $x_{d}$. In D3Net, we continuously feed the restored $I^{\prime}$ from the degradation decomposition branch back into the feature maps to enhance the restoration capability of the restoration reconstruction branch, i.e. $\hat{I}=f_{\text{rec}}(I,I^{\prime})$. This feedback mechanism encourages both branches to collaborate in jointly managing image restoration and correction, thereby enhancing the overall quality of image restoration.
@@
$\mathcal{L}_{1}=\|\hat{I}-I\|_{1},$ (1)
where $\hat{I}$ represents the reconstructed image and $I$ represents the ground truth image.
@@ III-B The Cross-Domain Degradation Analyzer
To address the interplay of these effects, CDDA employs Fourier transform analysis [53] to study the spectral characteristics of images. Specifically, the methodology proceeds as follows. First, for each image $I$ in a batch, we convert its color space and normalize it to prepare for frequency domain analysis. The two-dimensional discrete Fourier transform of image $I$ is represented as:
$F(u,v)=\sum_{x=0}^{M-1}\sum_{y=0}^{N-1}I(x,y)e^{-j2\pi\left(\frac{ux}{M}+\frac{vy}{N}\right)},$ (2)
@@
Here, $Q$ denotes the query, $K$ signifies the key, and $V$ represents the value, each originating from different feature sources, i.e., $D_{f}$ or $I$. In generating $C_{t}$, $I$ serves as both $Q$ and $K$, while $D_{f}$ acts as $V$, undergoing non-linear projection through $\phi(\cdot)$, which consists of three layers of $3\times 3$ convolutions. This design ensures that $C_{t}$ is tightly aligned with spatial image features, emphasizing the fine-grained degradation details that need correction, while injecting frequency-based degradation prompt from $D_{f}$ to enhance accuracy.
@@ III-C1 Decision Units
The goal of DUs is to enable the network to make decisions based on the input degradation features and the two-level prompt at the current processing stage. A DU consists of a multi-layer convolutional neural network structure, which can effectively compress the feature space through stacked convolutional layers and pooling layers, extracting independent hard decision vectors to ultimately determine whether to activate the corresponding ADB. Mathematically, the processing of a DU can be described as a function $\mathcal{F}_{DU}$, which maps input features to a decision space. For the input degraded feature map $x^{i}$ at the $i$-th stage, this mapping can be formalized as:
$h^{i}=\mathcal{F}^{i}_{DU}(x^{i};S_{t}),$ (8)
where $h^{i}=[h^{i}_{0},h^{i}_{1}]\in\mathbb{R}^{2}$ represents the output signal of the $i$-th DU, and $S_{t}$ represents the strategy prompt.
@@
$x_{d}^{i}=\frac{\exp((\log(h^{i})+g^{i})/\tau)}{\sum^{2}_{j=1}\exp((\log(h^{j})+g^{j})/\tau)},$ (9)
where $x_{d}^{i}$ is the output of the $i$-th decision unit, $g^{i}$ is noise sampled from the Gumbel distribution, and the temperature parameter $\tau$ is set to 1. This function approximates the sampling process of a discrete distribution while maintaining differentiability, facilitating the end-to-end training.
Importantly, the DUs and $S_{t}$ work together: while $S_{t}$ encodes a global restoration strategy derived from cross-domain interactions, the DUs use this strategic knowledge to adaptively choose which decomposition steps to apply, ensuring efficient and targeted feature refinement.
@@
$Y^{i}=Y^{i-1}+(x^{i}+\mathcal{F}_{ADB}^{i}(x_{c}^{i}))\cdot x_{d,1}^{i},\quad i=1,\ldots,N,$ (10)
where $Y^{i}$ is the output after the $i$-th stage, $x^{i}$ is the input to the $i$-th ADB, $\mathcal{F}_{ADB}^{i}(\cdot)$ denotes the $i$-th ADB’s operations, and $x_{c}^{i}$ represents the degraded features at the $i$-th stage. The factor $x_{d,1}^{i}$ is the activation probability obtained from the Gumbel-Softmax, serving as a smooth gating mechanism. $N$ is the number of DUs, which is set to 12 in this paper. At inference time, a hard decision can be made by replacing $x_{d,1}^{i}$ with the binary indicator $(x_{d,1}^{i}>\alpha)$, where $\alpha=0.5$. In this way, our DDM ensures stable gradient transmission during end-to-end training while achieving highly flexible and dynamic adjustments.
@@ III-C2 Adaptive Decomposition Blocks
Having briefly mentioned ADBs earlier, this section delves deeper into their operational mechanics, especially the adaptive decomposition of pixel-level degradation features through the utilization of the degradation correction prompt. To achieve this, we design a module that combines the input degradation feature map with the degradation correction prompt, which constitutes the basic component of the dynamic decomposition mechanism proposed in this paper. Specifically, each ADB consists of five sub-blocks: two convolutional blocks containing $g^{n}$onv [56], interleaved with two $1\times 1$ convolution blocks, and an output layer. Notably, we have improved $g^{n}$onv. Specifically, $C_{t}$ is integrated as a direct input to each $g^{n}$onv layer alongside the degradation feature $x_{g}$, ensuring that the decomposition operation is informed by explicit correction instructions at each pixel:
$\begin{split}[p_{0}^{HW\times C},q_{0}^{HW\times C}]=\phi_{\text{concat}}([{x_{g}},C_{t}])&\in\mathbb{R}^{HW\times 2C},\\ \end{split}$ (11)
where $x_{g}\in\mathbb{R}^{HW\times C}$ is the degraded features to be decomposed, and $\phi_{\text{concat}}$ concatenates these features and the degradation correction prompt $C_{t}\in\mathbb{R}^{HW\times C}$ along the channel dimension. This change not only enhances the information flow between channels but also allows the degradation correction prompt to directly participate in the adjustment of each feature. Next, a depthwise separable convolution $f$ is employed to model local spatial relationships and perform degradation corrections:
@@
${Y_{g}^{j}}=\phi_{\text{out}}({p}_{1})\in\mathbb{R}^{HW\times C},\quad j=1,\ldots,3.$ (14)
In summary, unlike traditional all-in-one methods that struggle to handle varying degradation types, D3Net separates the restoration process into degradation correction and restoration strategy guidance, using CDDA and DDM in tandem. This synergy enables D3Net to efficiently adapt to the nuances of degradation, improving the overall restoration quality without incurring the computational burden typically associated with handling multiple degradation modes. The integration of frequency-domain insights into the restoration process provides a deeper understanding of degradation patterns, allowing for accurate and targeted restoration. Moreover, D3Net’s unique approach to dynamic feature refinement, driven by guided prompts, ensures that restoration strategies are dynamically tailored to specific degradation characteristics, offering a high level of precision and effectiveness across a variety of image restoration tasks.

### v2

@@ Abstract
Currently, restoring clean images from a variety of degradation types using a single model is still a challenging task. Existing all-in-one image restoration approaches struggle with addressing complex and ambiguously defined degradation types. In this paper, we introduce a dynamic degradation decomposition network for all-in-one image restoration, named D3Net. D3Net achieves degradation-adaptive image restoration with guided prompt through cross-domain interaction and dynamic degradation decomposition. Concretely, in D3Net, the proposed Cross-Domain Degradation Analyzer (CDDA) engages in deep interaction between frequency domain degradation characteristics and spatial domain image features to identify and model variations of different degradation types on the image manifold, generating degradation correction prompt and strategy prompt, which guide the following decomposition process. Furthermore, the prompt-based Dynamic Decomposition Mechanism (DDM) encourages the network to adaptively select restoration strategies utilizing the two-level prompt generated by CDDA. Thanks to the synergistic cooperation between CDDA and DDM, D3Net achieves superior flexibility and scalability in handling unknown degradation, while effectively reducing unnecessary computational overhead. Extensive experiments on multiple image restoration tasks demonstrate that D3Net outperforms the state-of-the-art approaches, especially improving PSNR by 5.47dB on SOTS-Outdoor and 3.30dB on GoPro, respectively.
@@
Image restoration, prompt learning, dynamic network, low-level vision.
@@ I Introduction
Figure 1: Conceptual differences between our work and other existing all-in-one methods (better viewed by zooming in).
Image restoration aims to remove degradations from corrupted images and generate high-quality ones. It is widely applied in various real-world scenarios, such as autonomous driving, digital photography, and security surveillance. Currently, most image restoration approaches focus on overcoming a single type of degradations, which may be denoising [1, 2, 3, 4, 5], deblurring [6, 7, 8, 9, 10], deraining [11, 12, 13, 14, 15], dehazing [16, 17, 18, 19], or low-light enhancement [20, 21, 22, 23, 24, 25]. Although these previous research has yielded promising results in a specific application, the challenge of addressing multiple degradation tasks in image restoration with a single model still persists. This significantly hampers the practical applicability of the existing image restoration techniques.
@@
As shown in Fig. 1(d), in this paper, we propose the Dynamic Degradation Decomposition Network (D3Net) for all-in-one image restoration, which aims to address unknown multiple degradation situations. Unlike previous all-in-one image restoration approaches [26, 27, 28, 29, 30, 31], our method departs from the perspective of frequency domain and dynamic networks to effectively handle unknown and complex degradations without relying on explicit degradation labels or predefined prompts. In order to address the first question as above, we design a Cross-Domain Degradation Analyzer (CDDA) that generates degradation-related prompt to guide the degradation decomposition process. Specifically, it considers the complementary degradation characteristics between frequency and spatial domains, deeply interacting key degradation features in the frequency domain with spatial image features to identify and model the degradation types on the image manifold. Meanwhile, it generates degradation correction prompt and strategy prompt, thereby providing guidance for the following degradation decomposition process. Moreover, for the second question, we introduce a degradation-aware Dynamic Decomposition Mechanism (DDM). The design of DDM aims to progressively and dynamically decompose degradation features, fully leveraging the two-level prompt generated by CDDA to guide the image restoration process. Specifically, this mechanism distinctly enables the network to dynamically adapt the input from diverse decomposition blocks, tailoring its model to suit different degradation scenarios.
@@
  • We introduce a novel dynamic decomposition mechanism. Owing to its adaptive prompt-guided mode, this mechanism can dynamically select an appropriate restoration strategy based on the degradation state of the images, thereby achieving considerable degradation decomposition effects while reducing computational overhead.
@@ II-A Image Restoration
Image Restoration for Single Degradation. Traditional image restoration techniques mainly target at a specific type of degradation, such as blur, noise, or haze. These approaches [32, 33, 34, 35, 36], optimized for specific problems, can often achieve good results in the corresponding degradation scenario. For example, Mohan et al. [32] propose a model-based approach for addressing blur caused by camera shake. However, such approaches typically rely on specific prior degradation models, such as linear blur or motion blur. These model-based approaches may be severely limited when the actual degradation process does not match the assumed one. Recently, with the advancement of deep learning, numerous image restoration approaches based on deep learning have emerged and performed excellently across various image restoration tasks, such as denoising [1, 2, 3, 4, 5], deblurring [6, 7, 8, 9, 10], deraining [11, 12, 13, 14, 15, 37], dehazing [16, 17, 18, 19], and low-light enhancement [20, 21, 21, 22, 23, 24, 25]. However, constrained by model architectures designed for single tasks, they tend to perform exceptionally well in single degradation removal tasks but struggle to deal with other types of degradation.
@@
Similarly, Conde et al. [31] propose a method utilizing predefined degradation texts to guide multiple degradation types. However, this approach does not deeply model the inherent degradation features within the images. By focusing solely on user-provided textual instructions without in-depth analysis of degradation patterns in the images, it may overlook critical details necessary for optimal restoration, especially when degradations are subtle or intertwined.
@@ II-B Dynamic Networks
The research on dynamic networks has gradually garnered attention in the academic community, with the objective of reducing computational complexity while ensuring model performance. In recent years, researchers introduce various methodologies to achieve this goal. For example,  [40] introduces the groundbreaking BL-DNN approach, which concatenates two distinct deep CNNs with varying depths. When the softmax prediction accuracy of the first CNN surpasses a predefined threshold, the network terminates prematurely to reduce computational complexity. Subsequent works [41, 42, 43, 44] have also improved recognition tasks based on this concept. Moreover, some researchers develop methods that dynamically adjust convolutional kernel weights [45, 46, 47, 48] and kernel shapes [49, 50, 51] during inference based on their inputs. However, due to their irregular memory access and computation patterns, these methods are limited in further efficiency improvement. Diverging from most existing dynamic networks, we design a degradation prompt-based dynamic decomposition mechanism that dynamically selects appropriate image restoration strategies.
@@ III-A The Architecture of D3Net
In order to achieve prompt-guided adaptive image degradation restoration, we design a D3Net as illustrated in Fig. 2. It primarily consists of a restoration reconstruction branch and a degradation decomposition branch. The restoration reconstruction branch employs a generic U-Net structure [52] to reconstruct the input image. Building upon this foundational process, for a given degraded input image $I$, the shallow degradation features $x$ extracted by the first encoder of the restoration reconstruction branch are inputted into the degradation decomposition branch. That is, $x=f_{\text{encoder}}(I)$. Concurrently, $I$ is fed into the cross-domain degradation analyzer, which generates degradation correction prompt $C_{t}$ and strategy prompt $S_{t}$ to guide the degradation decomposition branch. Under the guidance of $C_{t}$ and $S_{t}$, it then iteratively decomposes the features through multiple stages. Specifically, at each stage $k$ ($k=1,2,...,N$), the branch selects degradation-adaptive strategies to decompose the current features $Y^{k-1}$ into refined degradation-decomposed $Y^{k}$ and an updated restoration strategy $\rho^{k}$, i.e.,
$Y^{k},\rho^{k}=f_{\text{ddm}}(Y^{k-1},C_{t},S_{t}).$
After $N$ stages of decomposition, the final degradation-decomposed features $Y^{N}$ are passed through a feature refinement module to generate the decomposition-finalized features $I^{\prime}$. In D3Net, we continuously feed $I^{\prime}$ from the degradation decomposition branch back into the feature maps to enhance the restoration capability of the restoration reconstruction branch, i.e. $\hat{I}=f_{\text{rec}}(I,I^{\prime})$. This feedback mechanism encourages both branches to collaborate in jointly managing image restoration and correction, thereby enhancing the overall quality of image restoration.
@@
$\mathcal{L}_{1}=\|\hat{I}-I_{gt}\|_{1},$ (1)
where $\hat{I}$ represents the reconstructed image and $I_{gt}$ represents the ground truth image.
@@ III-B The Cross-Domain Degradation Analyzer
To address the interplay of these effects, CDDA employs Fourier transform analysis [53] to study the spectral characteristics of images. Specifically, the methodology proceeds as follows: First, for each image $I$ in a batch, we convert its color space and normalize it to prepare for frequency domain analysis. The two-dimensional discrete Fourier transform of image $I$ is represented as:
$F(u,v)=\sum_{x=0}^{m-1}\sum_{y=0}^{n-1}I(x,y)\cdot e^{-j2\pi\left(\frac{ux}{m}+\frac{vy}{n}\right)},$ (2)
@@
Here, $Q$ denotes the query, $K$ signifies the key, and $V$ represents the value, each originating from different feature sources, i.e., $D_{f}$ or $I$. In generating $C_{t}$, $I$ serves as both $Q$ and $K$, while $D_{f}$ acts as $V$, undergoing non-linear projection through $\phi(\cdot)$, which consists of three layers of $3\times 3$ convolutions. This design ensures that $C_{t}$ is tightly aligned with spatial image features, emphasizing the fine-grained degradation details that need correction, while injecting frequency-based degradation prompt from $D_{f}$.
@@ III-C1 Decision Units
The goal of DUs is to enable the network to make decisions based on the input degradation features and the two-level prompt at the current processing stage. A DU consists of a multi-layer convolutional neural network structure, which can effectively compress the feature space through stacked convolutional layers and pooling layers, extracting independent hard decision vectors to ultimately determine whether to activate the corresponding ADB. Mathematically, the processing of a DU can be described as a function $\mathcal{F}_{DU}$, which maps input features to a decision space. For the input degraded feature map $Y^{i-1}$ at the $i$-th stage, with $Y^{0}$ being the initial degraded feature map, this mapping can be formalized as:
$h^{i}=\mathcal{F}^{i}_{DU}(Y^{i-1},S_{t}),$ (8)
where $h^{i}=[h^{i}_{0},h^{i}_{1}]\in\mathbb{R}^{2}$ represents the output signal of the $i$-th DU, with $h^{i}_{0}$ and $h^{i}_{1}$ denoting the logits for suppressing or activating the corresponding ADB, respectively. $S_{t}$ represents the strategy prompt.
@@
$\rho^{i}_{m}=\frac{\exp((h^{i}_{m}+g^{i}_{m})/\tau)}{\sum_{k=0}^{1}\exp((h^{i}_{k}+g^{i}_{k})/\tau)},\quad m=0,1,$ (9)
where $\rho^{i}_{m}$ denotes the probability of selecting option $m$ for the $i$-th decision unit, with $m=0$ for suppression and $m=1$ for activation. Here, $g^{i}_{m}$ is independent and identically distributed (i.i.d.) noise sampled from the Gumbel distribution for the $m$-th option, and $\tau$ is a temperature parameter controlling the smoothness of the approximation. This function approximates the sampling process of a discrete distribution while maintaining differentiability, facilitating end-to-end training.
Importantly, the DUs and $S_{t}$ work together: while $S_{t}$ encodes a global restoration strategy derived from cross-domain interactions, the DUs use this strategic knowledge to adaptively choose which decomposition stages to apply, ensuring efficient and targeted feature refinement.
@@
$Y^{i}=Y^{i-1}+\mathcal{F}_{ADB}^{i}(Y^{i-1})\odot\rho^{i}_{1},\quad i=1,\ldots,N,$ (10)
where $Y^{i}$ is the output after the $i$-th stage, $Y^{i-1}$ is the input feature from the previous stage, $\mathcal{F}_{ADB}^{i}(\cdot)$ denotes the $i$-th ADB’s operations, and $\rho^{i}_{1}$ is the activation probability obtained from the Gumbel-Softmax corresponding to $h^{i}_{1}$, serving as a differentiable gating mechanism that controls the contribution of the ADB’s output. We anneal the temperature $\tau$ from 1.0 to 0.1 over the training process, encouraging $\rho^{i}_{1}$ to approach a near-discrete distribution (close to 0 or 1) while remaining differentiable. At inference time, we compute $\rho^{i}_{1}$ using the deterministic Softmax formulation in Eq. 9 with $g^{i}_{m}=0$, ensuring stable decision-making without gradient computation. $N$ is the number of DUs, which is set to 12 in our work. In this way, our DDM ensures stable gradient transmission during end-to-end training while achieving highly flexible and dynamic adjustments.
@@ III-C2 Adaptive Decomposition Blocks
Having briefly mentioned ADBs earlier, this section delves deeper into their operational mechanics, especially the adaptive decomposition of pixel-level degradation features through the utilization of the degradation correction prompt. To achieve this, we design a module that combines the input degradation feature map with the degradation correction prompt, which constitutes the basic component of the dynamic decomposition mechanism proposed in this paper. Specifically, each ADB consists of five sub-blocks: two convolutional blocks containing $g^{n}$onv [56], interleaved with two $1\times 1$ convolution blocks, and an output layer. Notably, we have improved $g^{n}$onv. Specifically, $C_{t}$ is integrated as a direct input to each $g^{n}$onv layer alongside the degradation feature $Y_{g}$, ensuring that the decomposition operation is informed by explicit correction instructions at each pixel:
$\begin{split}[p_{0}^{HW\times C},q_{0}^{HW\times C}]=\phi_{\text{concat}}([{Y_{g}},C_{t}])&\in\mathbb{R}^{HW\times 2C},\\ \end{split}$ (11)
where $Y_{g}\in\mathbb{R}^{HW\times C}$ is the degraded features to be decomposed, and $\phi_{\text{concat}}$ concatenates these features and the degradation correction prompt $C_{t}\in\mathbb{R}^{HW\times C}$ along the channel dimension. This change not only enhances the information flow between channels but also allows the degradation correction prompt to directly participate in the adjustment of each feature. Next, a depthwise separable convolution $f$ is employed to model local spatial relationships and perform degradation corrections:
@@
${Y_{g}^{j}}=\phi_{\text{out}}({p}_{1})\in\mathbb{R}^{HW\times C},\quad j=1,2,3.$ (14)
Here, $j$ denotes the order of the processing using the $g^{n}$onv module at the current stage.
In summary, unlike traditional image restoration methods that struggle to handle multiple degradation types, D3Net separates the restoration process into degradation correction and restoration strategy guidance, using the proposed CDDA and DDM respectively. This synergy enables D3Net to efficiently adapt to the nuances of degradation, improving the overall restoration quality without incurring the computational burden. The integration of frequency-domain insights into the restoration process provides a deeper understanding of degradation patterns, allowing for accurate and targeted restoration. Moreover, D3Net’s unique approach to dynamic feature refinement, driven by guided prompts, ensures that restoration strategies are dynamically tailored to specific degradation characteristics, offering a high level of precision and effectiveness across a variety of image restoration tasks.